1) Determine the gametes produced by each genotype. Calculate the frequency of each gamete type. a. J c. JjKk d. jjkK e. CcDdEeFf f. DDeeGGhh
Answer #1

Answer a -Jj - if completely linked then n=1, if independent assortment then n=2, 2^2=4

Answer b -JJ-n=1,2^1=2

Answer C-JjKk, n=2, then 2^2=4

Answer D -CcDdEeFf- if completely linked then n=1, if independent assortment then n=4then 2^4=16

Answer E -jjKK-n=2, then 2^2=4

Answer F -DDeeGGhh-  n=4 then 2^4=16.